** INCLUDE MATLAB CODE **

Use MATLAB® to implement and plot the discrete fourier transform (DFT) for the following waveform: x(t) = 5 + 2cos(2*pi*t – pi/2) + 3cos(2*pi*4*t)
Answer #1

`Hey,

Note: In case of any queries, just comment in box I would be very happy to assist all your queries

clc
clear all
close all
t = 0:1/100:10-1/100; % Time vector
x = 5 + 2*cos(2*pi*t - pi/2) + 3*cos(2*pi*4*t); % Signal
y = fft(x); % Compute DFT of x
m = abs(y); % Magnitude
y(m<1e-6) = 0;
p = unwrap(angle(y)); % Phase
f = (0:length(y)-1)*100/length(y); % Frequency vector
subplot(2,1,1)
plot(f,m)
title('Magnitude')
subplot(2,1,2)
plot(f,p*180/pi)
title('Phase')

Kindly revert for any queries

Thanks.